OpenStudy (safia21):

Find the perimeter of a rectangle, given that its length is 14 cm and its area is 84 cm².

OpenStudy (lalaly):

84=14*W W=6 perimeter = 2*6 + 14*2=40

OpenStudy (anonymous):

A=l*w 84=14w w=6 P=2(l+w) P=2(14+6)=2(20)=40